# Splines-Based Feature Importance in Kolmogorov-Arnold Networks: A Framework for Supervised Tabular Data Dimensionality Reduction

### v1

@@ 4 Methodology
We propose a family of feature importance measures derived directly from Kolmogorov-Arnold Networks (KANs). Specifically, we consider four complementary approaches: (i) coefficient-based importance using $L^{1}$ and $L^{2}$ norms of spline weights (KAN-L1, KAN-L2), (ii) knock-out importance based on loss differentials (KAN-KO), and (iii) sensitivity-based importance computed from input gradients (KAN-SI).
@@ 4.1 Coefficient-Based Feature Importance (KAN-L1 and KAN-L2)
Recall from Eq.( 3) that each feature $x_{j}$ in a KAN layer is associated with a spline weight block
$W_{\text{spline}}^{(j)}\;\in\;\mathbb{R}^{m\times K},$ (9)
where $m$ is the number of outputs and $K$ the number of spline basis functions. This block collects all coefficients linking feature $j$ to every output unit.
The coefficient-based feature importance is then defined by the norm of $W_{\text{spline}}^{(j)}$:
$I_{L_{1}}(x_{j})\;=\;\|W_{\text{spline}}^{(j)}\|_{1}=\sum_{p=1}^{m}\sum_{k=1}^{K}|w_{pjk}|,\qquad I_{L_{2}}(x_{j})\;=\;\|W_{\text{spline}}^{(j)}\|_{2}=\left(\sum_{p=1}^{m}\sum_{k=1}^{K}w_{pjk}^{2}\right)^{1/2}.$ (10)
Finally, o enable comparability across features, these scores are normalized:
$\tilde{I}_{L_{1}}(x_{j})\;=\;\frac{I_{L_{1}}(x_{j})}{\sum_{i=1}^{d}I_{L_{1}}(x_{i})},\qquad\tilde{I}_{L_{2}}(x_{j})\;=\;\frac{I_{L_{2}}(x_{j})}{\sum_{i=1}^{d}I_{L_{2}}(x_{i})}.$ (11)
Extension to base and spline coefficients.
While the definitions above use only the spline weight matrices $W_{i}$, the same construction can be applied to the base weight matrix $W_{\text{base}}\in\mathbb{R}^{m\times d}$. In this case, the $i$-th column of $W_{\text{base}}$ represents the linear contribution of feature $x_{i}$, and its $L^{1}$ or $L^{2}$ norm can be taken as an additional importance score. Combining the norms of both $W_{\text{base}}$ and $W_{\text{spline}}$ yields a measure of the total contribution of each feature, whereas restricting to $W_{\text{spline}}$ isolates the nonlinear component of feature importance.
@@ 4.2 KAN-Knockout (KAN-KO)
Let the trained KAN layer (see Eq.( 3)) be parameterized by $W:=(W_{\text{base}},W_{\text{spline}})$. For a given feature index $j\in\{1,\dots,d\}$, we define the knockout operator $\mathcal{K}_{j}$ acting on the first KAN layer by
@@
$l^{(-j)}(X)=l(X)-\big[W_{\text{base}}(:,j)\,\phi(X_{j})+W_{\text{spline}}^{(j)}\,b_{j}(X_{j})\big].$ (13)
To quantify the contribution of each input feature, we evaluate how much the task loss increases when that feature is removed from the model. Let $\ell(\hat{Y},Y)$ denote the task loss and $P$ the data distribution over $(X,Y)$. We write $f_{W}$ for the full KAN model, where $W$ collects the base and spline weights (see  3.1). The expected or population risk of the model is
@@
$\tilde{I_{j}}\;=\;\frac{\Delta_{j}}{\sum_{k=1}^{d}\Delta_{k}+\delta},$ (17)
@@ 4.3 KAN- Sensitivity Integral (KAN-SI) Feature Importance
$I_{i}=\mathbb{E}_{\mathbf{x}\sim P}[S_{i}(\mathbf{x})],\hskip 14.22636pt\widehat{I_{i}}=\dfrac{1}{N}\sum_{n=1}^{N}\left|\frac{\partial f(\mathbf{x^{n}})}{\partial x_{i}}\right|$ (19)
where $\widehat{I}_{i}$ is the empirical estimator over a held-out set $\{\mathbf{x}^{n}\}_{n-1}^{N}$ (validation or out-of-fold) to avoid optimistic bias. From (1), let the inner pre-activations be
@@
and for B-splines of degree $p$, only the basis derivatives $\big(B^{\prime}_{k}(x_{i})\big)$ of at least degree $p+1$ are nonzero at a given $x_{i}$ (local support), making (20) efficient to evaluate.
Since sensitivities depend on the measurement of $x_{i}$, we report a normalized score.
@@
where $\mathrm{Std}$ (standart deviation) or $\mathrm{IQR}$ (interquartile range) is calculated on the same split used for (19). For one - hot encoded categoricals, compute $\tilde{I}$ per dummy and sum back to the original category.
@@ 4.4 Assessment Procedure (From Feature Selection to Prediction)
Let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\begin{cases}\{1,\dots,C\},&\text{classification},\\ \mathbb{R},&\text{regression}.\end{cases}$.
Let $\mathcal{S}=\{s_{1},\dots,s_{L=9}\}$ denote the feature selectors. Each $s\in\mathcal{S}$ fitted on training data returns a nonnegative importance vector $a_{s}\in\mathbb{R}_{+}^{d}$, normalized so $\sum_{j=1}^{d}a_{s,j}=1$. For a retention ratio $k\in\{20\%,40\%,60\%\}$, set $n_{k}=\max\!\left\{1,\,\left\lceil\frac{k\,\cdot\,d}{100}\right\rceil\right\},$ representing the number equivalent of k% of the features.
@@
For each task $t\in\{\text{classification},\text{regression}\}$, we consider a family of predictors $\mathcal{P}_{t}=\{P_{t,1},\dots,P_{t,M=4}\}$, including Logistic Regression (LogReg) / Ridge, Random Forests (RF), Gradient Boosted Trees (GBT), and XGBoost (XGB). Predictive performance is quantified by

### v2

@@ 4 Methodology
In a KAN, each input coordinate is passed through a small set of one-dimensional spline functions and then combined linearly, so all parameters associated with a given feature form an explicit, low-dimensional block. This structure naturally supports several ways of quantifying how much each input dimension contributes to the learned mapping. We therefore define four KAN-based feature-importance criteria derived from a trained model. This section discusses our four KAN-based selectors.  [^3] See the appendix A for more details
@@ 4.1 Coefficient-based feature importance (KAN-L1 and KAN-L2)
In a KAN layer, each input feature $x_{i}$ is expanded into a set of $K$ B-spline basis functions,
$z_{i}=\sum_{k=1}^{K}w_{ik}\,B_{k}(x_{i}),$ (8)
where $w_{ik}$ are the learned spline coefficients for feature $x_{i}$. These coefficients capture how strongly the model relies on different local regions of $x_{i}$’s domain.
To quantify feature importance for $x_{i}$, we aggregate each coefficient vector $\mathbf{w}_{i}=(w_{i1},\dots,w_{iK})$ using its $\ell_{1}$ and $\ell_{2}$ norms:
$I_{L_{1}}(x_{i})=\|\mathbf{w}_{i}\|_{1}=\sum_{k=1}^{K}|w_{ik}|,\qquad I_{L_{2}}(x_{i})=\|\mathbf{w}_{i}\|_{2}=\left(\sum_{k=1}^{K}w_{ik}^{2}\right)^{1/2}.$ (9)
To enable comparability across features, we normalize these scores to create the final importance scores:
$\tilde{I}_{L_{1}}(x_{i})=\frac{I_{L_{1}}(x_{i})}{\sum_{j=1}^{d}I_{L_{1}}(x_{j})},\qquad\tilde{I}_{L_{2}}(x_{i})=\frac{I_{L_{2}}(x_{i})}{\sum_{j=1}^{d}I_{L_{2}}(x_{j})}.$ (10)
These normalized quantities provide a direct and interpretable measure of how much each input contributes, on average, through its spline expansion.
@@ 4.2 KAN-Knockout (KAN-KO)
Let the trained KAN layer (see Subsection 3.1) be parameterized by $W:=(W_{\text{base}},W_{\text{spline}})$. For a given feature index $j\in\{1,\dots,d\}$, define the knockout operator $\mathcal{K}_{j}$ acting on the first KAN layer by
@@
$y^{(-j)}(x)=y(x)-W_{\text{base}}(:,j)\,\phi(x_{j})-W_{\text{spline}}^{(j)}\,b_{j}(x_{j}).$ (12)
To quantify the contribution of each input feature, we evaluate how much the task loss increases when that feature is removed from the model. Let $\ell(\hat{y},y)$ denote the task loss and $P$ the data distribution over $(x,y)$. We write $f_{W}$ for the full KAN model, where $W$ collects the base and spline weights (see Subsection 3.1). The expected or population risk of the model is
@@
$I_{j}=\frac{\Delta_{j}}{\sum_{k=1}^{d}\Delta_{k}+\delta},$ (16)
@@ 4.3 KAN- Sensitivity Integral (KAN-SI) Feature importance
$I_{i}=\mathbb{E}_{\mathbf{x}}[S_{i}(\mathbf{x}],\hskip 14.22636pt\widehat{I_{i}}=\dfrac{1}{N}\sum_{n=1}^{N}\left|\frac{\partial f(\mathbf{x^{n}})}{\partial x_{i}}\right|$ (18)
where $\widehat{I}_{i}$ is the empirical estimator over a held-out set $\{\mathbf{x}^{n}\}_{n-1}^{N}$ (validation or out-of-fold) to avoid optimistic bias.
From (1), let the inner pre-activations be
@@
and for B-splines of degree $p$, only $p+1$ basis derivatives $B^{\prime}_{k}(x_{i})$ are non-zero at a given $x_{i}$ (local support), which makes (19) efficient to evaluate.
Since the sensitivities depend on the measurement of $x_{i}$, we report a normalized score.
@@
where $\mathrm{Std}$(standart deviation) or $\mathrm{IQR}$(interquartile range) is computed on the same split used for (18). For one - hot encoded categoricals, compute $\tilde{I}$ per dummy and sum back to the original category. KAN-SI quantifies the importance of feature $x_{i}$ as the expected on-data magnitude of the model’s directional derivative given by (18) and report the scale-invariant score using (21) to neutralize unit effects.
@@ 4.4 Assessment Procedure (From Feature Selection to Prediction)
Let $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and
$y_{i}\in\begin{cases}\{1,\dots,C\},&\text{classification},\\ \mathbb{R},&\text{regression}.\end{cases}$ (22)
Let $\mathcal{S}=\{s_{1},\dots,s_{L}\}$ denote feature selectors (e.g., KAN-L1/L2, KAN-SI/KO, MI, LASSO, etc.). Each $s\in\mathcal{S}$ fitted on training data returns a nonnegative importance vector $a_{s}\in\mathbb{R}_{+}^{d}$, normalized so $\sum_{j=1}^{d}a_{s,j}=1$. For a retention ratio $k\in\{20\%,40\%,60\%\}$, set
$n_{k}=\max\!\left\{1,\,\left\lceil\frac{k\,\cdot\,d}{100}\right\rceil\right\},\text{number equivalent of k\% of the features }$ (23)
@@
For each task $t\in\{\text{classification},\text{regression}\}$, we consider a family of predictors $\mathcal{P}_{t}=\{P_{t,1},\dots,P_{t,M}\}$, including Logistic Regression / Ridge, Random Forests, Gradient Boosted Trees (GBT), and XGBoost. Predictive performance is quantified by
@@
We evaluate each selector-predictor pair using leakage-safe $F$-fold cross-validation. For each fold $f$, we fit preprocessing and the feature selector $s$ only on the training split $T_{f}$, obtain the selected feature set $J_{s,k}^{(f)}$, train the predictor on the projected training data, and then compute the score $S_{s,k,t,m,f}$ (macro-$F_{1}$ or $R^{2}$) on the projected validation split $V_{f}$. The overall performance $\mathrm{Score}(s,k,t,m)$ is the average of these fold-level scores. This procedure ensures that both feature selection and model training see only training data in each fold, preventing information leakage from the validation sets and yielding an unbiased estimate of generalization performance.  [^4] See the mathematical formulation of this process in appendix B.3
Comparison with Baseline Methods
We compare our spline-based feature selectors against Mutual Information (MI) ranking (Peng et al., 2005), Random Forest (Breiman, 2001), LASSO-based feature selection (Tibshirani, 1996b), and Recursive Feature Elimination (RFE) for Support Vector Machine (SVM) (SVM-RFE)(Guyon et al., 2002).
5 Experiments
Reproducibility details and data sets information can be found at Appendix B.1 and Appendix B.2.
The comparative analysis across datasets reveals that averaging F1 and $R^{2}$ scores across retention levels $k$ reveals consistent yet dataset-specific interactions between selectors, predictors, and the underlying data structure.
Classification datasets
On the Breast Cancer dataset (Figure 1), Gradient Boosted Trees perform best with KAN-L1, LASSO, and RF importance, often matching the full-feature baseline. In contrast, KAN-KO and Mutual Information yield poor subsets, while Logistic Regression and tree ensembles (RF, XGBoost) benefit most from sparsity- or impurity-based selectors. This indicates that selectors capturing either linear sparsity or tree-based splits are most effective, whereas knock-out and sensitivity KAN variants are less suited.
Figure 1: Average $F1$ Score ( averaged across 20/40/60% retained features) per Selectors for Each Classifiers
In the Digits dataset (Figure 1), the pattern shifts: Gradient Boosted Trees and XGBoost achieve higher F1 scores with LASSO, SVM-RFE, and several KAN variants (KAN-KO, KAN-SI, KAN-L1), sometimes surpassing the full-feature baseline. Here, dimensionality reduction removes redundancy and sharpens predictive power. By contrast, KAN-L2 and Mutual Information underperform, reflecting their limits in capturing complex feature interactions. For the synthetic make_classification dataset (Figure 1), all KAN-based selectors (KAN-L1/L2, KAN-SI, KAN-KO) consistently outperform classical baselines, confirming their advantage in structured, high-signal settings where ground-truth feature relevance extends beyond sparsity or univariate dependence. Finally, on the Wine dataset (Figure 1), the advantage shifts back toward classical selectors: RF importance, LASSO, Mutual Information, and SVM-RFE provide the strongest subsets, while most KAN-based selectors lag behind. Only KAN-SI remains competitive, suggesting that when features are few and relationships are well-structured, simpler selectors aligned with linear sparsity or impurity-based measures are more effective.
Regression datasets
On the California dataset (nonlinear, heterogeneous), Figure. 2, KAN-L2 and KAN-SI preserve tree-ensemble performance: with Random Forest and XGBoost they stay close to the orediction using all Features (minor drops $\sim$0.01-0.03), while Gradient Boosted Tree is slightly more sensitive.
Figure 2: Relative Average $R^{2}$ Score (averaged across 20/40/60% retained features) per Selectors for Each Regressors
In contrast, KAN-L1 is overly aggressive: it depresses Ridge markedly (large $R^{2}$ gap vs. All Features) and also trims XGBoost and Gradient Boosted Tree more than other KAN variants, indicating loss of weak but complementary signals. KAN-KO feature selector is generally competitive, trailing All Features by a small margin and often matching the classical Mitual Info, Random Forest and SVM-RFE subsets. For Diabetes dataset (small-$n$, noisy), Figure. 2), KAN-L1 again under-selects (lowest bars for Gradient Boosted Tree, Random Forest, XGBoost), whereas KAN-L2,KAN-SI yield the most stable KAN performance across predictors and typically track LASSO and Mutual Info. KAN-KO behaves as a conservative KAN subset, usually second-best among KANs. In the Diamond dataset (strong signal, structured, mixed types), Figure. 2, All models are high-$R^{2}$; here KAN-KO,KAN-L2 and KAN-SI are nearly indistinguishable from All Features for Random Forest, XGBoost and Gradient Boosted Tree and even outperform classical features selectors (Mitual Info, Random Forest and SVM-RFE subsets, LASSO and Random Forest). KAN-L1 is the only KAN variant that noticeably drops on some predictors, consistent with over-pruning informative but correlated attributes. For the Synthetic make regression dataset (well-conditioned, Figure. 2, All KAN variants cluster tightly with baselines across predictors (tiny spreads): when the signal is clean and features are already informative, KAN selection neither helps nor hurts much; KAN-L2/KAN-SI remain the safest choices.
Taken together, these findings suggest that KAN-based selectors are particularly attractive when feature interactions and nonlinearity play a central role, or when one wishes to couple selection and interpretability (see Appendix C). Their main failure mode is over-pruning correlated yet informative features under an $\ell_{1}$-style criterion (KAN-L1), in regression. More broadly, the results reinforce that dimensionality reduction is not uniformly harmful: when a selector is aligned with the structure of the data and the bias of the predictor, removing redundant or spurious variables can improve both accuracy and interpretability.
6 Conclusion
This work presented, to our knowledge, the first systematic study of Kolmogorov-Arnold Network (KAN) based feature selection on tabular classification and regression benchmarks. We defined four KAN-derived criteria (KAN-L1, KAN-L2, KAN-SI, KAN-KO) and compared them against widely used baselines, including LASSO, Random Forest feature importance, mutual information, and SVM-RFE, across multiple datasets and feature-retention levels. Empirically, KAN-L2, KAN-SI, and KAN-KO provide competitive and often superior performance in structured or strong-signal settings, while remaining robust on real-world tasks; KAN-L1 can be effective in classification but tends to over-prune in noisy or correlated regression problems. Classical selectors such as LASSO and Random Forest remain strong choices, particularly when they align with the assumptions of the downstream model. Furthermore, our stability and redundancy analyses show that KAN-based criteria produce reproducible feature subsets across folds and do not inflate correlation among selected features indicating that they offer both reliable and non-redundant selection even without enforcing sparsity constraints. Beyond aggregate scores, our spline-based case studies show that KANs can yield smooth, one-dimensional response functions that link feature values to model behaviour in a transparent way, offering an interpretable view of nonlinear and multivariate relevance that is difficult to obtain from purely sparsity- or impurity-based methods. Overall, our findings indicate that KAN-based feature selection is a practical, interpretable alternative to traditional approaches, and they provide guidance on when specific KAN criteria are most beneficial (e.g., KAN-L2/KAN-SI for noisy regression, KAN-SI/KAN-KO for interaction-heavy classification). We see this as a step toward feature-selection methods that jointly offer strong predictive performance, robustness to irrelevant variables, and clear mechanistic insight into how individual predictors influence model outputs. At the same time, training KANs is noticeably slower than fitting sparse linear models or tree ensembles, which limits their use in large-scale screening or AutoML settings. An important direction for future work is therefore to develop faster, yet still interpretable, KAN architectures and training schemes (e.g., lightweight KAN layers, pruning or distillation of spline components, or hybrid models that reuse KAN-based scores to guide simpler selectors).
Appendix
Part I
Appendix A Mathematical Motivation of KAN-Based Feature Scores
This section gives a brief functional motivation for the four KAN-based feature-importance criteria used in the main paper.
A.1 Functional view of the first KAN layer
For clarity, consider a single-output KAN with one hidden layer. The first layer applies, for each unit $r=1,\dots,R$ and feature $j=1,\dots,d$, a univariate spline
$g_{j,r}(x_{j})=\sum_{k=1}^{K}\theta_{j,r,k}\,B_{k}(x_{j}),$ (26)
where $\{B_{k}\}_{k=1}^{K}$ are fixed B-spline basis functions and $\theta_{j,r,k}$ are learned coefficients. Collecting all coefficients attached to feature $j$ in the first layer gives a parameter block
$\theta_{j}=(\theta_{j,r,k})_{r=1,\dots,R;\,k=1,\dots,K}.$ (27)
The network output can be written schematically as
$f(x)=\sum_{r=1}^{R}w_{r}\,\sigma\!\Big(\sum_{j=1}^{d}g_{j,r}(x_{j})\Big),$ (28)
so that feature $j$ influences $f$ only through its spline family $\{g_{j,r}\}_{r}$ controlled by $\theta_{j}$. The scores below quantify the importance of feature $j$ by measuring, in different ways, how much the learned function $f$ depends on this block.
A.2 Coefficient norms: KAN-L1 and KAN-L2
For fixed $j$ and $r$, define the spline $g_{j,r}(x_{j})=\sum_{k}\theta_{j,r,k}B_{k}(x_{j})$ and consider its $L^{2}(P_{X})$ norm with respect to the data distribution $P_{X}$:
$\|g_{j,r}\|_{L^{2}(P_{X})}^{2}=\mathbb{E}_{X}\!\Big[g_{j,r}(X_{j})^{2}\Big]=\theta_{j,r}^{\top}G_{j}\,\theta_{j,r},$ (29)
where $G_{j}=\mathbb{E}_{X}[B(X_{j})B(X_{j})^{\top}]$ is the Gram matrix of the B-spline basis for feature $j$. For standardized inputs and regular knots, $G_{j}$ is well-conditioned, so there exist constants $0<\lambda_{\min}\leq\lambda_{\max}<\infty$ such that
$\lambda_{\min}\,\|\theta_{j,r}\|_{2}^{2}\leq\|g_{j,r}\|_{L^{2}(P_{X})}^{2}\leq\lambda_{\max}\,\|\theta_{j,r}\|_{2}^{2}.$ (30)
Thus, up to fixed constants, the $\ell_{2}$ norm of the spline coefficients is proportional to the $L^{2}$ “energy” of the learned univariate transformation of feature $j$. Summing over hidden units $r$ yields that $\|\theta_{j}\|_{2}$ is a proxy for the total contribution of feature $j$ to the first-layer representation. This motivates the KAN-L2 score as
$\mathrm{score}_{\text{L2}}(j)=\|\theta_{j}\|_{2}.$ (31)
The KAN-L1 score $\mathrm{score}_{\text{L1}}(j)=\|\theta_{j}\|_{1}$ emphasizes sparsity over basis functions: features whose effect is concentrated on a small number of knots receive higher values, encouraging compact, interpretable univariate responses.
A.3 Sensitivity integral: KAN-SI
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be the trained KAN predictor and assume $f$ is differentiable almost everywhere. For a small perturbation $\delta$ along coordinate $j$ we have the first-order approximation
$f(x+\delta e_{j})-f(x)\approx\delta\,\frac{\partial f}{\partial x_{j}}(x),$ (32)
where $e_{j}$ is the $j$-th canonical basis vector. Taking expectations over the data distribution and random perturbations with $\mathbb{E}[|\delta|]=c$ gives
$\mathbb{E}\big[|f(X+\delta e_{j})-f(X)|\big]\approx c\,\mathbb{E}_{X}\!\Big[\big|\tfrac{\partial f}{\partial x_{j}}(X)\big|\Big].$ (33)
The quantity
$S_{j}^{\mathrm{SI}}=\mathbb{E}_{X}\!\Big[\big|\tfrac{\partial f}{\partial x_{j}}(X)\big|\Big]$ (34)
therefore measures the expected local sensitivity of the prediction to perturbations in feature $j$. Our KAN-SI score is the empirical estimate of $S_{j}^{\mathrm{SI}}$ computed on a validation set, using the gradient of the trained KAN. Features with large $S_{j}^{\mathrm{SI}}$ are those for which small changes in $x_{j}$ typically induce large changes in the output, making KAN-SI a derivative-based global importance measure.
A.4 Knock-out loss: KAN-KO
Let $\ell(f(X),Y)$ denote the training loss (cross-entropy for classification or squared error for regression) and $R(f)=\mathbb{E}[\ell(f(X),Y)]$ its risk. Given a trained KAN $f$, we define $f^{(-j)}$ as the same network but with all spline parameters associated with feature $j$ in the first layer set to zero, so that feature $j$ no longer contributes to the first-layer representation. The knock-out importance of feature $j$ is then
$\Delta_{j}=R(f^{(-j)})-R(f)\geq 0,$ (35)
the increase in risk incurred when the contribution of feature $j$ is ablated. In practice, we estimate $\Delta_{j}$ by the empirical risk difference on a held-out set, using the same trained parameters and only modifying the first-layer spline block of feature $j$. When $f$ approximately minimizes $R$, larger $\Delta_{j}$ indicate that the model relies heavily on feature $j$ for accurate predictions. This motivates the KAN-KO score as a leave-one-feature-out, loss-based measure of relevance.
A.5 Summary
In summary, the four KAN-based criteria exploit the spline-based parameterization of the first layer to approximate complementary notions of feature relevance: KAN-L2/KAN-L1 measure the “energy” and sparsity of the learned univariate transformations, KAN-SI captures gradient-based global sensitivity, and KAN-KO quantifies the increase in risk when a feature’s contribution is removed. All are computed directly from a trained KAN, without external wrapper search or additional surrogate models.
Appendix B Reproducibility
This section provides necessary reproducibility details
B.1 Reproducibility Details
[…]
B.2 Data Sets
To evaluate the effectiveness of the Kolmogorov-Arnold Networks dimensinality reduction methods, experiments were conducted on a diverse collection of benchmark datasets covering both classification and regression tasks. These datasets includes well-established repositories from scikit-learn as well as synthetically generated datasets and are summarized in Table 1.
Table 1: Benchmark and synthetic datasets
Task Dataset Observations Numbers of features References
Classification Wine 178 13 [Forina et al., 1991].
Breast Cancer Wisconsin 569 30 [Street et al., 1993]
Digits 11,797 64 [Pedregosa et al., 2011]
Make classification 500 10 [Pedregosa et al., 2011]
Regression California Housing 20,640 8 [Efron et al., 2004]
Diabetes 442 10 [Efron et al., 2004].
Diamonds 53,940 10 [Wickham, 2008]
Make regression 500 10 [Pedregosa et al., 2011]
B.3 Leakage-safe cross-validation